GFTOPK(1) GFTOPK(1) NAME gftopk-convertgenericfontfilestopackedfontfiles SYNOPSIS gftopk[-v]gffilename[pkfilename] DESCRIPTION Thegftopkprogramconvertsagenericfontfileoutputby,forexample,mf(1),toapackedfontfile forusebyDVI-readingprograms.Packedfontfiles(PKfiles)aremuchsmallerthanthecorre-spondingGFfiles,sotheyarerapidlybecomingthefontformatofchoice. Thegffilenameonthecommandlinemustbecomplete.Becausetheresolutionispartofthe extension,itwouldnotmakesensetoaddonadefaultextensionasisdonewithTeXorDVIreadingsoftware.Thepkfilenamedefaultstothesame(stripped)nameasgffilename,anditis placedinthecurrentworkingdirectorywiththe‘pk’suffixreplacing‘gf’.Forexample,theinput fileio.300gfwouldbecomeio.300pk. OPTIONS Unlessthe-vswitchisgiven,gftopkoperatessilently.With-v,theversionnumberoftheprogram andstatisticsaboutthepackingareoutputtostdout. ENVIRONMENT gftopklooksforgffilenamefirstinthecurrentdirectory.Ifitisnotpresentthere,itusesthe environmentvariableGFFONTS.Ifthatisnotset,itusesthevariableTEXFONTS.Ifthisis notset,itusesthesystemdefaults: TEXFONTS:/usr/TeX/lib/tex/fonts//:., Seetex(1)forthedetailsofthesearching. SEEALSO gftype(1),pktogf(1),pktype(1) DonaldKnuthetal.,METAFONTware AUTHORS TomasRokickiwrotetheprogram.ItwaspublishedaspartoftheMETAFONTwaretechnical report,availablefromtheTeXUsersGroup.PaulRichardsoriginallyportedittoUnix. 7Jan92 1